OpenStudy (anonymous):

Can you help me with these three questions? 1.Identify which equation is a quadratic equation. y = −6x − 3 y + x2 = (x − 1)(x + 7) y − 9 = (x − 1)(x2 + 4) y = 5x(x + 3) − 8 2.What are the solutions of 6x2 + 7x – 5 = 0? 3.Solve 3x2 = 12x.

OpenStudy (anonymous):

I think 3 is x = 0, x = 4 but I'm not sure.

OpenStudy (mandre):

Yes for 3 A quadratic equation is where the maximum power is 2. So simplify each equation to determine which is quadratic.

OpenStudy (anonymous):

^^^ Sorry I'm still not understanding what is a maximum power and where would that be? @Mandre Thanks!

OpenStudy (mandre):

In the following equation: x^4 + x^2 + x -20 the maximum power would be 4 because we have x^4 You understand?

OpenStudy (anonymous):

So would D, be the answer since the largest power is two?

OpenStudy (mandre):

What about B?

OpenStudy (anonymous):

C would be a maximum of three, right? So that couldn't be it.

OpenStudy (anonymous):

B could work because there is a maximum power of 2 but there are a lot to the power of two.

OpenStudy (mandre):

A and C is definitely not.

OpenStudy (anonymous):

I agree with that.

OpenStudy (mandre):

does not matter how many there are. You can add or subtract to make it one.

OpenStudy (mandre):

Can you pick more than one answer?

OpenStudy (anonymous):

No.

OpenStudy (mandre):

Then there is a problem. Both B and D are quadratic unless you typed it wrong.

OpenStudy (anonymous):

Thats how they should look.

OpenStudy (mandre):

Haha. Look at B again. Bring all terms to the left then you see what the problem is

OpenStudy (anonymous):

Yes, so what do you suggest I do. Lol

OpenStudy (mandre):

What happens to the x^2 if you bring all terms to the left once you've multiplied on the right?

OpenStudy (anonymous):

It would cancel out.

OpenStudy (anonymous):

Wait, no it won't let me try doing it again.

OpenStudy (anonymous):

I don't think I'm doing it right, everything is on one side equal to 0.

OpenStudy (mandre):

y + x2 = (x − 1)(x + 7) y + x2 = x2 +7x -x -7

OpenStudy (anonymous):

The x^2 would cancel out if I moved it over.

OpenStudy (mandre):

Yup. So problem solved right?

OpenStudy (anonymous):

So would D be the answer?

OpenStudy (mandre):

Yes.

OpenStudy (anonymous):

Okay that makes a lot more sense. Thank you for helping me you don't even know how much that helped me out I've been struggling with that problem for the past hour. Thanks for your time I really appreciate it!

OpenStudy (mandre):

My pleasure.